The sum of Nora's age and her grandmother's age is 71. Four times Nora's age is 6 less than her grandmothers age. Find their ages.

Answer by checkley75(3666)

N+G=71, N=71-G
4N=G-6
4(71-G)=G-6
284-4G=G-6
-4G-G=-6-284
-5G=-290
G=-290/-5
G=58 ANSWER FOR GRANDOTHER'S AGE
PROOF
N+58=71
N=71-58
N=13 AGE OF NORA.
4*13=58-6
52=52